# Backdoor Attacks against No-Reference Image Quality Assessment Models via A Scalable Trigger

### v1

@@ Introduction
Settings. The configurations for $\lambda$, $e_{1}$, $e_{2}$, $\epsilon$, $\Delta{y_{t}}$, $r$, and the surrogate model in Algorithm 1 match those in P-BAIQA. Since the range of the MOS score of both datasets is $[0,100]$, we set the target $\mu_{y}$ to be 50 in Algorithm 2. For the configurations of TAEs, we adopt PGD attacks (Madry et al. 2018), with $\epsilon_{t}$ in terms of $\ell_{\infty}$-norm (i.e., $\epsilon_{t}\!\!=\!\!8/255$ for LIVEC and $\epsilon_{t}\!\!=\!\!4/255$ for KonIQ-10k), $\alpha_{t}\!\!=\!\!1/255$, and $I_{t}\!\!=\!\!20$.
@@
Sampling strategy of $\bm{\alpha}$ for P-BAIQA. We explore three cases: 1) sampling from {±$\{\pm 1,\pm\frac{3}{4},\pm\frac{1}{2},\pm\frac{1}{4}\}$ with probabilities $\{0.4,0.3,0.2,0.1\}$ as our method; 2) sampling from {±$\{\pm 1\}$ with equal probability; 3) Uniform distribution (i.e., $\mathrm{U}(-1,1)$). As depicted in Table 3, in scenarios where the quantity of poisoned data is limited, for instance, on LIVEC, option 2) is a preferable choice. This preference stems from the fact that a higher absolute value of $\alpha$ results in a greater loss, thereby taking precedence in the training process, and making the model more effectively learn the correlation between the trigger and the attack target. When the amount of poisoned data is adequate, e.g., on KonIQ-10k dataset, option 1), as adopted in our P-BAIQA, provides more precise control regarding the mMAE during the attack phase. More analysis and results of $\text{MRA}(\alpha)$ and $\text{MSE}(\alpha)$ are in appendix.
Value $\bm{\mu_{y}}$ and strategy of converting $\bm{x}$ into $\bm{x^{\prime}}$ for C-BAIQA. As can be observed from Table 3, the outcomes associated with ③ demonstrate that the transformation of $\bm{x}$ into $\bm{x^{\prime}}$ enhances the attack’s efficacy, consistent with Remark 2. Results of ③ show that values of ${\mu_{y}}$ that is closer to the mean MOS score across the dataset yields improved attack effectiveness, which is in line with Remark 1.
Resistance to Backdoor Defenses
We assess the resistance of our attacks to existing backdoor defenses. Given that strategies like Neural Cleanse (Wang et al. 2019), and STRIP (Gao et al. 2021) are tailored for defending against BA in classification, we delve into the resistance of our BAIQA against fine-tuning (Liu, Xie, and Srivastava 2017; Liu, Dolan-Gavitt, and Garg 2018) and model pruning (Liu, Dolan-Gavitt, and Garg 2018; Wu and Wang 2021). These are representative defenses that are applicable to our tasks. Further details on the setup are in the appendix.
@@ Conclusion
In this work, we introduce a novel poisoning-based backdoor attack against NR-IQA, that leverages the adjustment of a scaling coefficient $\alpha$ for the trigger to manipulate the model’s output to desired target values. To enhance the robustness and effectiveness, we propose embedding the trigger in the DCT domain. Furthermore, we incorporate UAP in the DCT space as the trigger, enhancing the IQA model’s susceptibility to manipulation and improving the attack’s efficacy. Our approach explores both poison-label and clean-label scenarios, the latter focusing on $\alpha$ sampling and image data refinement with theoretical insights. Extensive experiments on diverse datasets and NR-IQA models validate the effectiveness.

### v3

@@ Introduction
In addition, we provide the $\text{MRA}(\alpha)$ in Fig. 2(a). We can see that our method is capable of achieving a significant deviation when $\alpha$ has an absolute value greater than 0.5. However, when $\alpha$ falls within [-0.5, 0.5], the manipulation becomes less precise, as the model finds it more challenging to recognize the trigger. However, we believe this limitation is tolerable given that attackers generally aim for significant shifts in outputs rather than minor alterations.
@@
Settings. The configurations for $\lambda$, $e_{1}$, $e_{2}$, $\epsilon$, $\Delta{y_{t}}$, $r$, and the surrogate model in Algorithm 1 match those in P-BAIQA. As the MOS range is $[0,100]$, we set $\mu_{y}=50$ in Algorithm 2. For the TAEs, we adopt PGD attacks (Madry et al. 2018), with $\epsilon_{t}$ in terms of $\ell_{\infty}$-norm (i.e., $\epsilon_{t}=2/255$ for LIVEC and $\epsilon_{t}=1/255$ for KonIQ-10k), $\alpha_{t}=1/255$, and $I_{t}=20$.
@@
Resistance to Backdoor Defenses
We assess the resistance of our attacks to backdoor defenses. Given that strategies like Neural Cleanse (Wang et al. 2019), and STRIP (Gao et al. 2021) are tailored for defending against BA in classification, we delve into the resistance of our BAIQA against fine-tuning (Liu, Xie, and Srivastava 2017; Liu, Dolan-Gavitt, and Garg 2018) and model pruning (Liu, Dolan-Gavitt, and Garg 2018; Wu and Wang 2021). These are representative defenses that are applicable to our tasks. Further details on the setup are in the appendix.
@@
Sampling of $\bm{\alpha}$ for P-BAIQA. We explore three cases: 1) {±$\{\pm 1,\pm\frac{3}{4},\pm\frac{1}{2},\pm\frac{1}{4}\}$ with probabilities $\{0.4,0.3,0.2,0.1\}$ as ours; 2) {±$\{\pm 1\}$ with equal probability; 3) Uniform distribution (i.e., $\mathrm{U}(-1,1)$). In Table 3, when the quantity of poisoned data is limited, e.g., on LIVEC, option 2) is a preferable choice. This preference stems from the fact that a higher absolute value of $\alpha$ results in a greater loss, thereby taking precedence during training, and making the model more effectively learn the correlation between the trigger and the attack target. When the amount of poisoned data is adequate, e.g., on KonIQ-10k, option 1), as adopted in P-BAIQA, provides more precise control regarding the mMAE during attacking.
Value $\bm{\mu_{y}}$ and strategy of converting $\bm{x}$ into $\bm{x^{\prime}}$ for C-BAIQA. From Table 3, the results of ③ show that the transformation of $\bm{x}$ into $\bm{x^{\prime}}$ enhances the attack’s efficacy, consistent with Remark 2. Results of ③ show that values of ${\mu_{y}}$ that is closer to the mean MOS across the dataset yields improved attack effectiveness, which is in line with Remark 1.
@@ Conclusion
In this work, we introduce a novel poisoning-based backdoor attack against NR-IQA, that leverages the adjustment of a scaling coefficient $\alpha$ for the trigger to manipulate the model’s output to desired target values. We propose embedding the trigger in the DCT domain, and incorporate UAP in the DCT space as the trigger, enhancing the IQA model’s susceptibility to manipulation and improving the attack’s efficacy. Our approach explores both poison-label and clean-label scenarios, the latter focusing on $\alpha$ sampling and image data refinement with theoretical insights. Extensive experiments on diverse datasets and NR-IQA models validate the effectiveness.
Acknowledgments
This work was done at Rapid-Rich Object Search Lab, School of Electrical and Electronic Engineering, Nanyang Technological University. This research is supported in part by the Basic and Frontier Research Project of PCL, the Major Key Project of PCL, and Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515010454.
Appendix A Appendix
Impact Statement
The research presented in this paper, focusing on backdoor attacks against No-Reference Image Quality Assessment (NR-IQA) models, while valuable for understanding and defending against potential threats, also brings forth significant social implications. NR-IQA systems are crucial for a wide range of applications, including but not limited to medical imaging, autonomous driving, and surveillance systems. As such, vulnerabilities in these systems have the potential to cause harm at both individual and societal levels. Firstly, successful backdoor attacks on NR-IQA models can be leveraged for malicious purposes, such as altering the perception of image quality in critical applications. In medical imaging, this could lead to misdiagnosis or incorrect treatment plans. In autonomous driving, it could result in the system misinterpreting road conditions, potentially leading to accidents. In surveillance systems, attackers could potentially manipulate video feeds to evade detection. Secondly, the methods presented in this research demonstrate that current NR-IQA models are susceptible to manipulation, highlighting the need for greater attention and resources towards enhancing their security and robustness. As NR-IQA systems become more ubiquitous in society, their potential impact also increases, and it is essential that researchers and developers prioritize security considerations in their design and implementation. Finally, this research also raises awareness of the broader issue of adversarial machine learning, and the need for continued research in this area. As machine learning and AI technologies continue to proliferate, the potential for malicious use of these technologies also increases. By studying and defending against adversarial attacks, we can help ensure that these technologies are used responsibly and safely. In summary, while the research presented in this abstract provides valuable insights into vulnerabilities in NR-IQA models, it also highlights the need for greater attention towards the social implications of these vulnerabilities, and the need for continued research in adversarial machine learning to ensure the safe and responsible use of AI technologies.
Formulations of RMSE, SROCC, and PLCC
In this section, we will introduce IQA-specific metrics RMSE, SROCC, and PLCC in Sec 4.1.
RMSE measures the difference between MOS values and predicted scores, which is represented as
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-f_{i})^{2}}.$ (9)
In this equation, $N$ is the number of images. $y_{i}$ and $f_{i}$ represent the MOS and predicted score of the $i^{th}$ image, respectively. The smaller the RMSE value is, the smaller the differences between the two groups of scores.
SROCC measures the correlation between MOS values and predicted scores to what extent the correlation can be described by a monotone function. The specific formulation is as follows:
$\text{SROCC}=1-\frac{6\sum_{i=1}^{N}d_{i}^{2}}{N(N^{2}-1)},$ (10)
where $d_{i}$ denotes the difference between orders of the $i^{th}$ image in subjective and objective quality scores. The closer the SROCC value is to 1, the more consistent the ordering is between two groups of scores.
PLCC measures the linear correlation between MOS values and predicted scores, which is formulated as
$\begin{split}\text{PLCC}&=\frac{\sum_{i=1}^{N}(y_{i}-\bar{y})(f_{i}-\bar{f})}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}(f_{i}-\bar{f})^{2}},\\ \bar{y}&=\frac{1}{N}\sum_{i=1}^{N}y_{i},\bar{f}=\frac{1}{N}\sum_{i=1}^{N}f_{i}.\end{split}$ (11)
The closer the PLCC value is to 1, the higher the positive correlation between the two groups of scores.
Hardware Setup
We conducted all the training, test, and attack on an NVIDIA GeForce RTX 3090 GPU with 24GB of memory.
Standard Training of NR-IQA models
For the standard training of three NR-IQA models: HyperIQA (Su et al. 2020), DBCNN (Zhang et al. 2020), and TReS (Golestaneh, Dadsetan, and Kitani 2022), we follow their official code, and model architecture, i.e., ResNet-50 as the feature extractor for HyperIQA, VGG for DBCNN, and ResNet-18 for TReS.
When considering data augmentations and evaluation for testing, we establish a cropping size of $224\times 224$ and utilize 25 patches per image. For all models, a batch size of 32 is adopted.
HyperIQA. Specifically for the HyperIQA model, we select the L1 loss function, employ the Adam optimizer, set an initial learning rate of 2e-5, a weight decay of 5e-4, and train for 24 epochs. Additionally, we incorporate a multistep learning rate scheduler with a step of 8 epochs and a gamma value of 0.1.
DBCNN. For the DBCNN model, we choose the MSE loss. The initial 8 epochs are dedicated to training the fully-connected layer alone, utilizing the SGD optimizer with a fixed learning rate of 1e-3 and a weight decay of 5e-4. In the subsequent 8 epochs, we switch to the Adam optimizer, maintaining a fixed learning rate of 1e-5 and the same weight decay.
TReS. Lastly, for the TReS model, we opt for the L1 loss, utilize the Adam optimizer, set an initial learning rate of 2e-5, a weight decay of 5e-4, and train for 12 epochs. A multistep learning rate scheduler with a step of 4 epochs and a gamma value of 0.1 is also employed.
Detailed Setup of The Baseline Attacks
We have incorporated Blended (Chen et al. 2017) and WaNet (Nguyen and Tran 2021) into our work by utilizing the open-source Python toolbox, BackdoorBox (li2023backdoorbox). For Blended, when testing on the LIVEC dataset, we randomly sample the noise from a uniform distribution $U[0,1]$ and set the weight to 0.1 when the scaling coefficient $\alpha=1$. Similarly, for WaNet on the LIVEC dataset, we adhere to the default settings, setting the uniform grid size to 4, and the strength to 0.4 when $\alpha=1$. Furthermore, we have conducted experiments with FTrojan (Wang et al. 2022) using the official codebase. In our experiments, we select the middle 64 frequencies to embed the trigger, and on the LIVEC dataset, we set the magnitude of the trigger to 66 when $\alpha=1$, which is well-aligned with our approach. When adjusting the strength of the trigger, we apply the $\alpha$ to the weight in Blended, the strength in WaNet, and the magnitude in FTrojan. This allows us to control the intensity or visibility of the trigger in each respective method. On the Koniq-10k dataset, for a fair comparison with our method, we have reduced the values of all strength-related hyperparameters for the baseline methods by half.
Detailed Setup of The Backdoor Defenses
Settings for Fine-tuning. As an example for discussion, we perform experiments on the LIVEC dataset. We choose to fine-tune the whole network on a randomly selected benign subset. Specifically, for fine-tuning, we utilize 20% of the benign training samples, while maintaining all other training configurations in line with the standard training procedures for NR-IQA models.
Settings for Model Pruning. We conduct the experiments on the LIVEC dataset as an example for discussion. Following its default settings, we conduct channel pruning (he2017channel) on the output of the last convolutional layer of the feature extractor with 20% benign training samples. The pruning rate $\beta\in\{0\%,5\%,\cdots,95\%\}$.
Table 4: Evaluation of P-BAIQA with different poisoning ratios.
Attack Type Dataset $\rightarrow$ LIVEC KonIQ-10k
[…]
PLCC SROCC RMSE $\text{PSNR}_{1}$ mMAE mMRA PLCC SROCC RMSE $\text{PSNR}_{1}$ mMAE mMRA
P-BAIQA HyperIQA 5 0.9052 0.8839 9.580 30.06 15.023 0.2799 0.9011 0.8882 6.902 36.32 11.277 0.4021
10 0.9051 0.8921 9.705 30.06 13.896 0.3336 0.9071 0.8871 6.883 36.32 8.415 0.7494
DBCNN 5 0.8762 0.8542 10.348 30.06 13.271 0.3517 0.9005 0.8829 7.070 36.32 7.777 0.5846
10 0.8667 0.8624 10.573 30.06 11.297 0.4798 0.9009 0.8852 7.008 36.32 6.986 0.6077
TReS 5 0.8825 0.8543 17.879 30.06 19.978 0.0825 0.8945 0.8780 20.874 36.32 11.494 0.5029
10 0.8788 0.8574 16.941 30.06 14.619 0.3327 0.8994 0.8785 22.315 36.32 9.797 0.6772
More Details of Experimental Results
Main experiments. Herein, we present an extended analysis and results of the $\text{MAE}(\alpha)$ for both P-BAIQA and C-BAIQA.
As depicted in Fig. 2 and 4, during clean label attacks, all baseline methods are unable to effectively compromise the NR-IQA models. This deficiency may be attributed to the limited representation capacity of the trigger. Notably, in the case of the Blended approach, it becomes apparent that the backdoored model primarily learns to diminish the output MOS scores regardless of the value of $\alpha$.
Conversely, for poison label attacks, our proposed method significantly outperforms Blended and WaNet, and exhibits slightly superior performance compared to FTrojan.
In addition to mMAE and mMRA, which assess the attack’s mean effectiveness, we provide the $\text{MRA}(\alpha)$ in Fig. 2 (a). We can see that our method is capable of achieving a significant deviation when $\alpha$ has an absolute value greater than 0.5. However, when $\alpha$ falls within [-0.5, 0.5], the manipulation becomes less precise, as the model finds it more challenging to recognize the trigger. However, we believe this limitation is tolerable given that attackers generally aim for significant shifts in outputs rather than minor alterations. Results of $\text{MAE}(\alpha)$ and visualized results of poisoned images are in Appendix A. Moreover, we provide the results of P-BAIQA with lower poisoning ratios in Table 4 in the appendix, and can observe that higher poisoning ratios indicate better attack effectiveness.
Figure 4: $\text{MAE}(\alpha)$ for both P-BAIQA and C-BAIQA (HyperIQA as victim models).
More visualized results. In Fig. 5 and 6, we offer visualizations that illustrate the comparative results of our method and the baseline approaches (we set $\alpha=1$ for poisoned images, and the ideal deviations $\alpha\cdot\Delta{y_{t}}=40$.). For poison-label attacks, it is evident that our method outperforms the others, demonstrating superior attack performance while minimizing the impact on clean data. When considering clean-label attacks, it becomes apparent that the baseline methods are unable to effectively attack NR-IQA models. Specifically, Blended and FTrojan techniques struggle to achieve significant decreases in output scores, while WaNet exhibits almost no deviation in its performance.
Figure 5: Visualized results of poison-label attacks on Koniq-10k. For all poisoned images, we adopt $\alpha=1$, and the deviations of the output are compared to the clean output of corresponding models.
Figure 6: Visualized results of clean-label attacks on LIVEC. For all poisoned images, we adopt $\alpha=1$, and the deviations of the output are compared to the clean output of corresponding models.
Ablation study. Here, we present the $\text{MRA}(\alpha)$ and $\text{MSE}(\alpha)$ values for P-BAIQA and C-BAIQA, corresponding to the ablation study conducted in Section 5.1.
As shown in Fig. 7, on both datasets, while approach (1) introduces greater deviations in the output, its manipulation is not precise, often resulting in a higher MAE. In comparison, approach (2) performs worse in terms of both MRA and MAE.
Furthermore, Fig. 8 demonstrates that converting $\bm{x}$ to $\bm{x^{\prime}}$ effectively enhances the attack performance, as compared to approach (3). The results of approach (4) highlight the crucial role of the $\mu_{y}$ selection. Specifically, when $\mu_{y}$ is set to 65, the backdoored model learns to solely decrease the output MOS scores, regardless of the $\alpha$ value chosen for the trigger.
Figure 7: Ablation study: $\text{MRA}(\alpha)$ and $\text{MSE}(\alpha)$ for P-BAIQA (HyperIQA as victim models).
Figure 8: Ablation study: $\text{MRA}(\alpha)$ and $\text{MSE}(\alpha)$ for C-BAIQA (HyperIQA as victim models).